OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = -2x - 8 and contains the point (4, 4).

OpenStudy (anonymous):

Given equation of y=-2x-8. Gradient of this line is -2. Gradient of line that is perpendicular to given line = 1/2 Substitute point (4,4) and m=1/2 to the general equation of a line y=mx+c 4 = (1/2) (4) +c c= 2 Eqn. of line, y=(1/2)x +2

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

Np.